### v2

# On the Eldan–Gross inequality

Abstract.
A recent discovery of Eldan and Gross states that there exists a universal $C>0$ such that for all Boolean functions $f:\{-1,1\}^{n}\to\{-1,1\}$,
$\intop\nolimits_{\{-1,1\}^{n}}\sqrt{s_{f}(x)}d\mu(x)\geq C\textnormal{Var}(f)\sqrt{\log\left(1+\frac{1}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}$
where $s_{f}(x)$ is the sensitivity of $f$ at $x$, $\textnormal{Var}(f)$ is the variance of $f$, $\textnormal{Inf}_{j}(f)$ is the influence of $f$ along the $j$-th variable, and $\mu$ is the uniform probability measure. In this note, we give an alternative proof that applies to biased discrete hypercube, and spaces having positive Ricci curvature lower bounds in the sense of Bakry and Émery.
2020 Mathematics Subject Classification. 68R05; 47D07.
Key words. Boolean functions. Isoperimetric inequality. Hypercontractivity. Bakry–Émery curvature-dimension condition.
1. Introduction
[…]
$D_{j}f(x):=\frac{f(x)-f(x^{\oplus j})}{2},\qquad 1\leq j\leq n,$
where $x^{\oplus j}$ means flipping the $j$-th variable of $x$. Put
$|\nabla f|:=\left(\sumop\displaylimits_{j=1}^{n}|D_{j}f|^{2}\right)^{1/2}.$
For Boolean functions $f:\{-1,1\}^{n}\to\{-1,1\}$ the square of discrete gradient $|\nabla f(x)|^{2}$ coincides with the sensitivity of $f$ at $x$:
$|\nabla f(x)|^{2}=s_{f}(x):=\sharp\{1\leq j\leq n:f(x)\neq f(x^{\oplus j})\}.$
Recall that the influence $\textnormal{Inf}_{j}(f)$ of a Boolean function $f:\{-1,1\}^{n}\to\{-1,1\}$ along the $j$-th variable is the probability of “the value of $f$ changes when flipping $x_{j}$", or equivalently,
$\textnormal{Inf}_{j}(f)=\mathbb{E}|f-\mathbb{E}_{x_{j}}f|^{p},$
where $\mathbb{E}$ is defined with respect to the uniform distribution $\mu_{n}$, and $p>0$ is arbitrary. As usual, $\textnormal{Var}(f)$ is understood as the variance of $f$ with respect to the uniform distribution:
$\textnormal{Var}(f)=\mathbb{E}|f-\mathbb{E}f|^{2}=\mathbb{E}f^{2}-(\mathbb{E}f)^{2}.$
Note that for Boolean functions $f:\{-1,1\}^{n}\to\{-1,1\}$:
$\|f-\mathbb{E}f\|_{1}=4a(1-a)=\textnormal{Var}(f)\qquad\textnormal{ with }\qquad a=\mu_{n}(\{f=1\}).$ (1)
Here and in what follows $\|f\|_{p}$ denotes the $L^{p}$-norm with respect to the uniform probability measure unless otherwise stated.
Recently, Eldan and Gross [EG22] proved the following inequality.
Theorem 1 ([EG22]).
There exists a universal constant $C>0$ such that for all $n\geq 1$ and all Boolean function $f:\{-1,1\}^{n}\to\{-1,1\}$ we have
$\mathbb{E}|\nabla f|\geq C\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (2)
This inequality was motivated by a conjecture of Talagrand [Tal97] asked on the $p$-biased hypercube, i.e.,
$\mathbb{E}\sqrt{h_{f}}\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)},$ (3)
where $h_{f}$ is supported on a set $A=\{x:f(x)=1\}$, and at each point $x\in A$, the value $h_{f}$ counts the number of edges joining $x$ with $A^{c}$, i.e., complement of $A$. Clearly we always have $|\nabla f|=\sqrt{h_{f}+h_{-f}}\geq\sqrt{h_{f}}$, therefore, the conjecture (3), if correct, would imply Eldan–Gross inequality (2). However, in general $\mathbb{E}|\nabla f|$ and $\mathbb{E}\sqrt{h_{f}}$ may behave quite differently as $n\to\infty$, for example, if $A=\{(1,\ldots,1)\}$ is a single point then $\mathbb{E}|\nabla f|=\frac{n+\sqrt{n}}{2^{n}}\gg\frac{\sqrt{n}}{2^{n}}=\mathbb{E}\sqrt{h_{f}}$.
The proof of Eldan and Gross of the inequality (2) uses stochastic analysis and also leads to other impressive consequences. There are now several other proofs of (2); see for example van Handel [Ros20], Eldan–Kindler–Lifshitz–Minzer [EKLM23] and Beltran–Ivanisvili–Madrid [BIM23, Remark 2]. The main aim of this paper is to illustrate that some key ingredients of these proofs extend to more general spaces and the Boolean cube structure is not essential. Let us remark that the proof in [Ros20] can achieve the same goal after some modifications. Here, we highlight the use of an isoperimetric-type inequality by Bobkov and Götze [BG99] in 1999, the way of using which can be of independent interest.
We shall prove Theorem 1 on the $p$-biased hypercubes; see Section 2 for details. The proof relies on Bobkov and Götze’s isoperimetric-type inequality [BG99] and the hypercontractivity.
It might be easier to understand the proof in the continuous setting, and one typical example is the Gauss space. In fact, one may go further. Let $(E,\mu,\Gamma)$ be a full Markov triple with $(P_{t})=(e^{-tL})$ being the associated diffuse Markov semigroup in the sense of [BGL14]. Here $E$ is a measure space with $\mu$ being some probability measure, and $\Gamma$ is the carré du champs operator associated to $(P_{t})=(e^{-tL})$:
$\Gamma(f)=fL(f)-\frac{1}{2}L(f^{2}),$
for suitable $f:E\to\real$. By “suitable" we refer to a certain class of functions for which $\Gamma(f)$ is well-defined. In the sequel, we omit the rigorous definition to keep presentation compact, and we refer to [BGL14, Chapter 3] for detailed definitions for $(E,\mu,\Gamma)$ being a full Markov triple. One crucial property here is the diffusion property, which may be understood as the chain rule for $L$, so that many functional inequalities follow from Bakry–Émery criterion
$\Gamma(P_{t}f)\leq e^{-2Kt}P_{t}\Gamma f$
for all $f$ and $t\geq 0$. In this case, we say that the Ricci curvature is bounded from below by $K$, or it satisfies Bakry–Émery curvature-dimension condition $\textnormal{CD}(K,\infty)$. Examples of such include the heat semigroups on Riemannian manifolds and $K$ can be chosen to be the lower bound of Ricci curvature, which motivated the name in the above general framework. We will be interested in the case when the Ricci curvature lower bound $K$ is positive, for which a typical example is the Gauss space with $(P_{t})$ being the Ornstein–Uhlenbeck semigroup. In this case, the curvature lower bound is $K=1$.
For any $n\geq 1$, denote $E^{n}$ the $n$-fold product of $(E,\mu)$ equipped with the product probability measure $\mu_{n}.$ We consider the tensor product of $(P_{t})$ over $E^{n}$, which will be denoted by $(P_{t})=(e^{-tL})$ again whenever no confusion can occur. For $f:E^{n}\to\real$, we define its variance
$\textnormal{Var}(f):=\mathbb{E}|f-\mathbb{E}f|^{2}=\mathbb{E}(f^{2})-(\mathbb{E}f)^{2},$
with respect to $\mu_{n}$. Again, for Boolean functions $f:E^{n}\to\{-1,1\}$ one has
$\|f-\mathbb{E}f\|_{1}=4a(1-a)=\textnormal{Var}(f)\qquad\textnormal{ with }\qquad a=\mu_{n}(\{f=1\}).$ (4)
Here and in what follows, for $f:E^{n}\to\real$, the $L^{p}$ norm $\|f\|_{p}$ is defined with respect to $\mu_{n}$. We define
$|\nabla f|:=\Gamma(f)^{1/2}$
which is compatible with the discrete hypercube case by computing $\Gamma(f)$ for the heat semigroup. For any Boolean function $f:E^{n}\to\{-1,1\}$ and $1\leq j\leq n$, we define the influence of $j$-th variable as
$\textnormal{Inf}_{j}(f):=\mathbb{E}|f-\mathbb{E}_{x_{j}}(f)|.$
There are different definitions of influences in this context; our definition (also known as $L^{1}$-influence or geometric influence) agrees with the one in discrete hypercube case explained above.
Theorem 2.
Let $(E,\mu,\Gamma)$ be a full Markov triple satisfying Bakry–Émery curvature-dimension condition $\textnormal{CD}(K,\infty)$ with $\mu$ a probability measure and $0<K<\infty$. Then there exist constants $C=C_{K},C^{\prime}=C_{K}^{\prime}$ depending only on $K$ such that for all $n\geq 1$ and all Boolean functions $f:E^{n}\to\{-1,1\}$
$\mathbb{E}|\nabla f|\geq C_{K}\textnormal{Var}(f)\sqrt{\log\left(\frac{e}{\textnormal{Var}(f)}\right)}$ (5)
and
$\mathbb{E}|\nabla f|\geq C_{K}^{\prime}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (6)
The inequality (5) was inspired by Talagrand’s inequality [Tal93]. Although this inequality can be directly derived from the stronger, so-called local Bobkov inequality well-known in the literature, we will provide an "alternative" proof in Section 3. This proof, while slightly longer, incorporates the local Bobkov inequality as one of its key components. This approach might seem excessive, but it is intentional, as the proof of (6) will closely follow the structure of the "alternative" proof of (5), and additioanly, the inequality (5) itself will be used in our proof of (6).
Acknowledgment
P.I. is supported by National Science Foundation CAREER-DMS-2152401. H.Z. is supported by National Science Foundation DMS-2453408.
2. Boolean cube case
In this section, we give an alternative proof of Theorem 1. The proof also extends to the biased case which we now explain. For any $p\in(0,1)$ we equip $\{-1,1\}$ with the probability measure $\mu_{p}=p\delta_{1}+(1-p)\delta_{-1}$. We use $\|f\|_{q}$ to denote the $L^{q}$-norm of $f:\{-1,1\}^{n}\to\real$ with respect to $\mu_{p}^{\otimes n}$. We write $\mathbb{E}=\mathbb{E}_{\mu_{p}^{\otimes n}}$ for short, and consider
$P_{t}=(e^{-t}\textnormal{id}+(1-e^{-t})\mathbb{E})^{\otimes n},\qquad t\geq 0$ (7)
which is a semigroup of unital positive linear operators over $\real{}^{\{-1,1\}^{n}}$ such that $P_{0}(f)=f$ and $\lim_{t\to\infty}P_{t}(f)=\mathbb{E}f$. Accordingly, we write $\textnormal{Var}(f)$ for $\textnormal{Var}_{\mu_{p}}(f)$ with respect to this biased $\mu_{p}$.
For any $f:\{-1,1\}^{n}\to\{-1,1\}$ and $1\leq j\leq n$, we define the $j$-influence of $f$ as
$\textnormal{Inf}_{j}(f):=\|f-\mathbb{E}_{x_{j}\sim\mu_{p}}f\|_{1}.$ (8)
In a literature of Boolean functions, in general the quantity $\|D_{j}f\|_{1}$ is defined as the ($L^{1}$) influence on the biased hypercube. Notice that, for $f:\{-1,1\}^{n}\to\{-1,1\}$ one has
$\|f-\mathbb{E}_{x_{j}\sim\mu_{p}}f\|_{1}=4p(1-p)\|D_{j}f\|_{1},$
and the constant factor of 4p(1p)4p(1-p) will make no essential difference for our discussion. So we keep our choice (8). We still put
$|\nabla f(x)|:=\left(\sumop\displaylimits_{j=1}^{n}|D_{j}f(x)|^{2}\right)^{1/2}.$
Then the main result of this section reads as follows.
Theorem 3.
Fix $p\in(0,1)$ and let the notation be as above. Then there exists a constant $C_{p}>0$ such that for all $n\geq 1$ and all Boolean functions $f:\{-1,1\}^{n}\to\{-1,1\}$ we have
$\mathbb{E}|\nabla f|\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (9)
One ingredient of the proof can be derived from an isoperimetric-type inequality of Bobkov and Götze [BG99]. The proof is inspired by the unbiased case [KOS08]. More discussions can be found in Appendix B. Similar estimates were obtained in [Ros20, Claim 8, items 1-2] via a more direct argument.
Lemma 4.
Let $p\in(0,1).$ Suppose that $t\geq 0$ and $f:\{-1,1\}^{n}\to\{-1,1\}$. Then
$\textnormal{Var}(f)-\textnormal{Var}(P_{t}f)=1-\mathbb{E}|P_{t}f|^{2}\leq 2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}\mathbb{E}|\nabla f|.$ (10)
In particular, sending $t\to\infty$, we have
$\|f-\mathbb{E}f\|_{1}=\textnormal{Var}(f)\leq 2\max\{p,1-p\}\sqrt{\pi}\mathbb{E}|\nabla f|.$ (11)
Proof.
According to [BG99, Theorem 2.3], for probability spaces $(\Omega_{i},\nu_{i}),1\leq i\leq n$ and their product $(\Omega,\nu)$, we have
$\mathcal{I}(\mathbb{E}_{\nu}h)\leq\sqrt{2}\mathbb{E}_{\nu}\sqrt{\sumop\displaylimits_{i}\textnormal{Var}_{\nu_{i}}(h)}$ (12)
for any measurable $h:\Omega\to\{0,1\}$, where $\mathcal{I}$ is the Gaussian isoperimetric profile $\mathcal{I}:[0,1]\to[0,1/\sqrt{2\pi}]$ given by $\mathcal{I}=\varphi\circ\Phi^{-1}$ with $\Phi(r)=\intop\nolimits_{-\infty}^{r}\varphi(x)dx$ and $\varphi(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$.
Now we choose $\Omega_{i}=\{-1,1\}$. The probability measure will be chosen as follows. Fix $x\in\{-1,1\}^{n}$ and $t\geq 0$, and consider the following linear functional $\varphi_{i}$
$\varphi_{i}(g):=P_{t}(g)(x_{i})=(1-e^{-t})[pg(1)+(1-p)g(-1)]+e^{-t}g(x_{i}),\qquad g:\Omega_{i}=\{-1,1\}\to\real,$
where $P_{t}$ is exactly the depolarizing semigroup (7) in dimension $n=1$. Then $\varphi_{i}$ is positive and $\|\varphi_{i}\|=\varphi_{i}(1)=1$. By Riesz representation theorem, there exists a probability measure $\nu_{i}$ over $\{-1,1\}$ such that
$\mathbb{E}_{\nu_{i}}(g)=\intop\nolimits_{\{-1,1\}}gd\nu_{i}=P_{t}(g)(x_{i}),\qquad g:\{-1,1\}\to\real.$
Therefore, in general dimension $n\geq 1$ one has
$\mathbb{E}_{\nu}(g)=P_{t}(g)(x),\qquad g:\{-1,1\}^{n}\to\real$ (13)
for $P_{t}$ in (7). Moreover, for $h:\{-1,1\}^{n}\to\{0,1\}$ we have
$\mathbb{E}_{\nu_{i}}(h(y))$ $=(1-e^{-t})[ph(y_{1},\dots,1,\dots,y_{n})+(1-p)h(y_{1},\dots,-1,\dots,y_{n})]+e^{-t}h(y_{1},\dots,x_{i},\dots,y_{n})$
$=(1-e^{-t})(pa+qb)+e^{-t}c$
where $a=h(y_{1},\dots,1,\dots,y_{n}),b=h(y_{1},\dots,-1,\dots,y_{n})$ and $c=h(y_{1},\dots,x_{i},\dots,y_{n})$. Here we also write q=1pq=1-p and let us denote $r:=\max\{p,q\}$. Thus using $a=a^{2},b=b^{2},c=c^{2}$ we obtain by simple computations: if $c=a$, then
$\textnormal{Var}_{\nu_{i}}(h(y))=\mathbb{E}_{\nu_{i}}(h(y))-\mathbb{E}_{\nu_{i}}(h(y))^{2}=q(1-e^{-t})(p+qe^{-t})(a-b)^{2}\leq r^{2}(1-e^{-2t})(a-b)^{2},$ (14)
while if $c=b$, then
$\textnormal{Var}_{\nu_{i}}(h(y))=\mathbb{E}_{\nu_{i}}(h(y))-\mathbb{E}_{\nu_{i}}(h(y))^{2}=p(1-e^{-t})(q+pe^{-t})(a-b)^{2}\leq r^{2}(1-e^{-2t})(a-b)^{2}.$ (15)
So in either case, we have
$\textnormal{Var}_{\nu_{i}}(h(x))\leq r^{2}(1-e^{-2t})|h(x_{1},\dots,x_{i},\dots,x_{n})-h(x_{1},\dots,-x_{i},\dots,x_{n})|^{2}=4r^{2}(1-e^{-2t})|D_{i}h(x)|^{2}$ (16)
and thus for all $h:\{-1,1\}^{n}\to\{0,1\}$
$\sumop\displaylimits_{i=1}^{n}\textnormal{Var}_{\nu_{i}}(h(x))\leq 4r^{2}(1-e^{-2t})\sumop\displaylimits_{i=1}^{n}|D_{i}f(x)|^{2}=4r^{2}(1-e^{-2t})|\nabla h(x)|^{2}.$ (17)
Combining (12), (13) and (17), one has the pointwise inequality
$\mathcal{I}(P_{t}h)\leq 2\sqrt{2}\max\{p,1-p\}\sqrt{1-e^{-2t}}P_{t}|\nabla h|,\qquad h:\{-1,1\}^{n}\to\{0,1\}.$ (18)
For any $f:\{-1,1\}^{n}\to\{-1,1\}$ we set $h:=\frac{1}{2}(1+f)$ that takes values in $\{0,1\}.$ Then we just proved
$\mathcal{I}\left(\frac{1+P_{t}f}{2}\right)\leq\sqrt{2}\max\{p,1-p\}\sqrt{1-e^{-2t}}P_{t}|\nabla f|,\qquad f:\{-1,1\}^{n}\to\{-1,1\}.$ (19)
This, together with the the elementary inequality
$\mathcal{I}(s)\geq\sqrt{\frac{2}{\pi}}\left[\frac{1}{2}-2\left(\frac{1}{2}-s\right)^{2}\right],\qquad s\in[0,1]$ (20)
yields
$1-(P_{t}f)^{2}\leq 2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}P_{t}|\nabla f|,\qquad f:\{-1,1\}^{n}\to\{-1,1\}.$
Taking the expectation gives the desired estimate. ∎
Another ingredient is the hypercontractivity on the biased hypercube.
Proposition 5.
For $p\in(0,1)$ we consider the semigroup $P_{t}$ defined in (7). Then for any $1<r<2$ and for any $f:\{-1,1\}^{n}\to\real$ we have
$\|P_{t}f\|_{2}\leq\|f\|_{r}\qquad\textnormal{ for }\qquad t\geq 2p(1-p)\log\frac{1}{r-1}.$ (21)
When $p=1/2$, this is a well-known result by Bonami [Bon70] and the time in (21) is optimal. In the general biased case, the optimal time is more involved [Ole03]. We give a proof of (21) in the Appendix A for convenience.
From the above proposition, one may deduce the following lemma whose proof is essentially the same as that of [Ros20, Lemma 11].
Lemma 6.
Fix $p\in(0,1)$. For all $t\geq 0$ and for all $f:\{-1,1\}^{n}\to\real$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$ (22)
where $\theta(t)=\frac{1-e^{-2Kt}}{1+e^{-2Kt}}$ with $K=\frac{1}{4p(1-p)}$.
Proof.
For any $f$, consider ($\mathbb{E}_{x_{j}}:=\mathbb{E}_{x_{j}\sim\mu_{p}}$)
$f^{(0)}=f,\qquad f^{(j)}=\mathbb{E}_{x_{1},\dots,x_{j}}f,\quad 1\leq j\leq n.$
Write $f_{t}:=P_{t}(f)$. Then we have the martingale property
$\textnormal{Var}(f_{t})=\|f_{t}^{(0)}-f_{t}^{(n)}\|^{2}_{2}=\sumop\displaylimits_{j=1}^{n}\|f_{t}^{(j-1)}-f_{t}^{(j)}\|_{2}^{2}.$ (23)
Note that $\mathbb{E}_{x_{j}}P_{t}=P_{t}\mathbb{E}_{x_{j}}$, so by (23), hypercontractivity (21) and Hölder’s inequality
$\textnormal{Var}(f_{t})=\sumop\displaylimits_{j=1}^{n}\|f_{t}^{(j-1)}-f_{t}^{(j)}\|_{2}^{2}$ $=\sumop\displaylimits_{j=1}^{n}\|P_{t}(f^{(j-1)}-f^{(j)})\|_{2}^{2}$
$\leq\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1+e^{-2Kt}}^{2}$
$\leq\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2\theta(t)}\|f^{(j-1)}-f^{(j)}\|_{2}^{2-2\theta(t)}$
$\leq\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{2}^{2}\right)^{1-\theta(t)}$
$=\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$
where in the last equality we used (23) again. Since
$\|f^{(j-1)}-f^{(j)}\|_{1}=\|\mathbb{E}_{x_{1},\dots,x_{j-1}}(f-\mathbb{E}_{x_{j}}f)\|_{1}\leq\mathbb{E}|f-\mathbb{E}_{x_{j}}f|=\textnormal{Inf}_{j}(f),$
we get
$\textnormal{Var}(f_{t})\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)}.\qed$
Now we recall the following inequality due to [Tal93, Theorem 1.1], which will directly follow from a stronger local Bobkov inequality. In the next section one may find an alternative proof of it, though presented in the continuous setting.
Proposition 7.
Fix $p\in(0,1)$. Then there exists a universal constant $C_{p}>0$ such that for all $n\geq 1$ and all Boolean function $f:\{-1,1\}^{n}\to\{-1,1\}$ we have
$\mathbb{E}|\nabla f|\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(\frac{e}{\textnormal{Var}(f)}\right)}.$ (24)
Proof.
It is equivalent to proving the inequality for $h:\{-1,1\}^{n}\to\{0,1\}$. Starting from (18), we obtain the isoperimetric inequality by sending $t\to\infty$:
$\mathcal{I}(\mathbb{E}h)\leq 2\sqrt{2}\max\{p,1-p\}\mathbb{E}|\nabla h|,\qquad h:\{-1,1\}^{n}\to\{0,1\}$ (25)
Now instead of (20), we employ the following better lower bound
$\mathcal{I}(x)\geq Cx(1-x)\sqrt{\log\frac{1}{x(1-x)}},\qquad x\in(0,1).$ (26)
that entails
$\mathcal{I}(\mathbb{E}h)\geq C\textnormal{Var}(h)\sqrt{\log\frac{1}{\textnormal{Var}(h)}},\qquad h:\{-1,1\}^{n}\to\{0,1\}$ (27)
since $\textnormal{Var}(h)=\mathbb{E}h(1-\mathbb{E}h)$. Combining (25) and (27) finishes the proof. To see (26), it suffices to consider the cases when $x\downarrow 0$ and $x\uparrow 1$. Note that $\mathcal{I}(x)=\mathcal{I}(1-x)$, so it suffices to show (26) for $x\downarrow 0$ and this is known [Bob97, Remark 1]. ∎
Now we are ready to prove Theorem 3.
Proof of Theorem 3.
Denote $W(f):=\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}$. Recalling (11), the desired inequality (9) follows immediately when $W(f)$ is large enough, say $W(f)\geq 1/100$. So let us assume $W(f)<1/100$.
In view of (24), we may assume $\textnormal{Var}(f)\geq W(f)^{\alpha}$ with $\alpha>0$. In fact, if $\textnormal{Var}(f)<W(f)^{\alpha}$, then our desired (2) will follow from (24) because (recalling $W(f)<1/100$)
$\log\left(\frac{e}{\textnormal{Var}(f)}\right)=1+\log\left(\frac{1}{\textnormal{Var}(f)}\right)>1+\alpha\log\left(\frac{1}{W(f)}\right)\geq C\min\{1,\alpha\}\log\left(1+\frac{e}{W(f)}\right).$
We may also assume $\textnormal{Var}(f)\geq 100W(f)$. Otherwise, combining $\textnormal{Var}(f)<100W(f)$ and (24), to prove the desired inequality (2) it suffices to show
$\log\left(\frac{e}{100W(f)}\right)=\log\left(\frac{1}{100}\right)+\log\left(\frac{e}{W(f)}\right)\geq C\log\left(1+\frac{e}{W(f)}\right)$ (28)
which is true with universal $C>0$ since $W(f)<1/100$.
So from now on we assume $\textnormal{Var}(f)\geq\sqrt{W(f)}$ and $\textnormal{Var}(f)\geq 100W(f)$. According to Lemmas 4 and 6,
$\mathbb{E}|\nabla f|\geq\frac{\textnormal{Var}(f)-\textnormal{Var}(P_{t}f)}{2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}}\geq\frac{\textnormal{Var}(f)-W(f)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)}}{2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}}.$ (29)
where $\theta(t)=\frac{1-e^{-2Kt}}{1+e^{-2Kt}}$ with $K=\frac{1}{4p(1-p)}\geq 1$. Thus, to prove (9) it remains to show
$\frac{1-\left(\frac{W(f)}{\textnormal{Var}(f)}\right)^{\theta(t)}}{\sqrt{1-e^{-2Kt}}}\geq C\sqrt{\log\left(1+\frac{e}{W(f)}\right)}$ (30)
for some $t>0$. Writing $\epsilon=1-e^{-2Kt}$, (30) becomes
$\frac{1-\left(\frac{\textnormal{Var}(f)}{W(f)}\right)^{\frac{\epsilon}{\epsilon-2}}}{\sqrt{\epsilon}}\geq C\sqrt{\log\left(1+\frac{e}{W(f)}\right)}$ (31)
for some $\epsilon\in(0,1)$. Recalling $\textnormal{Var}(f)\geq 100W(f)$, we may choose $\epsilon^{-1}=\log\left(\textnormal{Var}(f)/W(f)\right)\in(1,\infty)$, so that
$\frac{1-\left(\frac{\textnormal{Var}(f)}{W(f)}\right)^{\frac{\epsilon}{\epsilon-2}}}{\sqrt{\epsilon}}=\frac{1-e^{\frac{1}{\epsilon-2}}}{\sqrt{\epsilon}}\geq C\sqrt{1+\frac{1}{\epsilon}}=C\sqrt{1+\log\left(\textnormal{Var}(f)/W(f)\right)},$
where we used the fact that
$\min_{\epsilon\in[0,1]}\frac{1-e^{\frac{1}{\epsilon-2}}}{\sqrt{1+\epsilon}}\geq\frac{1-e^{-1/2}}{\sqrt{2}}>0.$ (32)
Recall that $\textnormal{Var}(f)\geq\sqrt{W(f)}$. Therefore,
$\sqrt{1+\log\left(\textnormal{Var}(f)/W(f)\right)}\geq\sqrt{1+\frac{1}{2}\log\left(1/W(f)\right)}\geq C\sqrt{\log\left(1+\frac{e}{W(f)}\right)}.$ (33)
This concludes the proof of (31). ∎
3. Continuous analogues
In this section we prove Theorem 2. The main condition is that the Markov triple $(E,\mu,\Gamma)$ satisfies the Bakry–Émery curvature-dimension condition $\textnormal{CD}(K,\infty)$ with $\mu$ a probability measure and $K>0$. We refer to [BGL14] for more details.
We collect here several useful corollaries of $\textnormal{CD}(K,\infty)$ with $K>0$.
Proposition 8.
Fix a Markov triple $(E,\mu,\Gamma)$ with the diffuse Markov semigroup $P_{t}=e^{-tL}$. Assume that it satisfies curvature-dimension condition $\textnormal{CD}(K,\infty)$ with $K>0$.
  1. (1)
    Hypercontractivity [BGL14, Theorem 5.2.3, Proposition 5.7.1]: for $f:E\to\real$, $t\geq 0$ and $1<p\leq q<\infty$ we have
    $\|P_{t}f\|_{q}\leq\|f\|_{p}\qquad\textnormal{ for all }\qquad t\geq\frac{1}{2K}\log\frac{q-1}{p-1}.$ (34)
  2. (2)
    (Local) Bobkov inequality [BGL14, Theorem 8.5.3]: for $g:E\to\{0,1\}$ and $t\geq 0$ we have under suitable approximation
    $\mathcal{I}(P_{t}g)\leq\sqrt{\frac{1-e^{-2Kt}}{K}}P_{t}\sqrt{\Gamma(g)},$ (35)
    where $\mathcal{I}$ is the Gaussian isoperimetric profile.
Here, (35) is a special case of [BGL14, Theorem 8.5.3] that suffices for our use. In fact, the original [BGL14, Theorem 8.5.3] states that for $\alpha\geq 0,t\geq 0$ and measuable functions $f:E\to\real$
$\sqrt{\mathcal{I}^{2}(P_{t}f)+\alpha\Gamma(P_{t}f)}\leq P_{t}\left(\sqrt{\mathcal{I}^{2}(f)+c_{\alpha}(t)\Gamma(f)}\right),\qquad c_{\alpha}(t)=\frac{1-e^{-2Kt}}{K}+\alpha e^{-2Kt}.$ (36)
Then (35) follows by taking $\alpha=0$ in (36) and approximating the characteristic function $g:E\to\{0,1\}$ in (35) with measurable functions $f$ in (36). Here, the Boolean property plays a role, and we used $\mathcal{I}(0)=\mathcal{I}(1)=0$. For $n\geq 1$, we shall still denote by $P_{t}$ its $n$-fold products on $E^{n}$. Then the above results actually hold for $E^{n}$ after a tensorization argument, which we will use directly.
The key to the proof of Lemma 4 is (18) that can be considered as a discrete analog of (35). So following the same arguments in the proof Lemma 4, we may deduce
$1-(P_{t}f)^{2}\leq\sqrt{\frac{\pi(1-e^{-2Kt})}{2K}}P_{t}|\nabla f|,\qquad f:E^{n}\to\{-1,1\}$
from (35). Taking the expectation, we get
$\textnormal{Var}(f)-\textnormal{Var}(P_{t}f)=1-\mathbb{E}|P_{t}f|^{2}\leq\sqrt{\frac{\pi(1-e^{-2Kt})}{2K}}\mathbb{E}|\nabla f|.$ (37)
Sending $t\to\infty$ yields
$\|f-\mathbb{E}f\|_{1}=\textnormal{Var}(f)=1-\mathbb{E}|P_{t}f|^{2}\leq\sqrt{\frac{\pi}{2K}}\mathbb{E}|\nabla f|.$ (38)
Lemma 9.
Under the above assumptions, for $f:E^{n}\to\real$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta_{K}(t)}\textnormal{Var}(f)^{1-\theta_{K}(t)},$ (39)
for all $t\geq 0$ with $\theta_{K}(t)=\frac{1-e^{-2Kt}}{1+e^{-2Kt}}$.
Using the hypercontractivity estimate (34), the proof of the above lemma is the same as that of Lemma 6.
Proof of Theorem 2.
We first show (5), which may follow from the same proof of Proposition 7. Here we present another proof and the argument is similar to the proof of Theorem 3. When $\textnormal{Var}(f)\geq 1/100$, then the desired (5) follows immediately from (38).
Now we assume $\textnormal{Var}(f)<1/100$. By hypercontractivity (34) and Hölder’s inequality
$\textnormal{Var}(P_{t}f)=\|P_{t}(f-\mathbb{E}f)\|_{2}^{2}\leq\|f-\mathbb{E}f\|_{1+e^{-2Kt}}^{2}\leq\|f-\mathbb{E}f\|_{1}^{2\theta_{K}(t)}\|f-\mathbb{E}f\|_{2}^{2-2\theta_{K}(t)}=\textnormal{Var}(f)^{1+\theta_{K}(t)}.$ (40)
In the last equality we used the fact
$\|f-\mathbb{E}f\|_{1}=\|f-\mathbb{E}f\|_{2}^{2}=\textnormal{Var}(f)\qquad\textnormal{ for }\qquad f:E^{n}\to\{-1,1\}.$
This, together with (37), yields
$\mathbb{E}|\nabla f|\geq\sqrt{\frac{2K}{\pi}}\frac{\textnormal{Var}(f)-\textnormal{Var}(P_{t}f)}{\sqrt{1-e^{-2Kt}}}\geq\sqrt{\frac{2K}{\pi}}\frac{\textnormal{Var}(f)-\textnormal{Var}(f)^{1+\theta_{K}(t)}}{\sqrt{1-e^{-2Kt}}},\qquad t\geq 0.$ (41)
Thus, to prove (5), it suffices to show
$\frac{1-\textnormal{Var}(f)^{\theta_{K}(t)}}{\sqrt{1-e^{-2Kt}}}\geq C\sqrt{\log\left(\frac{e}{\textnormal{Var}(f)}\right)}$ (42)
for some $t>0$. Write $\epsilon=1-e^{-2Kt}$, and (42) becomes
$\frac{1-e^{\frac{\epsilon}{\epsilon-2}\log\left(\frac{1}{\textnormal{Var}(f)}\right)}}{\sqrt{\epsilon}}\geq C\sqrt{\log\left(\frac{e}{\textnormal{Var}(f)}\right)}$ (43)
for some $\epsilon\in(0,1)$. Recall that $\textnormal{Var}(f)<1/100$, so we may choose $\epsilon^{-1}=\log\left(\frac{1}{\textnormal{Var}(f)}\right)\in(1,\infty)$ and the proof is reduced to $\min_{\epsilon\in[0,1]}\frac{1-e^{\frac{1}{\epsilon-2}}}{\sqrt{1+\epsilon}}>0$. This is true as we argued in (32).
Therefore, we complete the proof of (5). Combining this with (38) and Lemma 9, one may establish (6) following exactly the same proof of Theorem 1 in the last section. ∎
Appendix A Proof of Proposition 5
Recall that $p\in(0,1)$ and $\mu_{p}(\{1\})=p,\mu_{p}(\{-1\})=1-p$. The generator of $P_{t}$ in dimension one is given by $L(f)=\mu_{p}(f)-f$ for $f:\{-1,1\}\to\real$. Put q:=1pq:=1-p. We shall prove the following logarithmic Sobolev inequality: for all non-negative $f:\{-1,1\}\to\real$ we have
$\mu_{p}(f^{2}\log f^{2})-\mu_{p}(f^{2})\log\mu_{p}(f^{2})\leq-\frac{1}{2pq}\mu_{p}(fL(f)).$ (44)
Then by standard arguments [Gro75], it implies
$\|P_{t}(f)\|_{r_{2}}\leq\|f\|_{r_{1}},\qquad t\geq 2pq\log\frac{r_{2}-1}{r_{1}-1}$ (45)
for all $f:\{-1,1\}^{n}\to\real$ and for all $1<r_{1}<r_{2}<\infty$ which finishes the proof of Proposition 5.
Now it remains to prove (44). By homogeneity, we may assume that $f(x)=1+sx$ with $s\in[-1,1]$. Then (44) becomes
$p(1+s)^{2}\log(1+s)^{2}+q(1-s)^{2}\log(1-s)^{2}-[p(1+s)^{2}+q(1-s)^{2}]\log[p(1+s)^{2}+q(1-s)^{2}]\leq 2s^{2},$
for $s\in[-1,1]$ by direct computations. Consider the function
$\psi(s):=p(1+s)^{2}\log(1+s)^{2}+q(1-s)^{2}\log(1-s)^{2}-[p(1+s)^{2}+q(1-s)^{2}]\log[p(1+s)^{2}+q(1-s)^{2}]-2s^{2}.$
Note that $\psi(0)=0$. It suffices to show $\psi(s)\leq\psi(0),s\in[-1,1]$. For this we compute
$\psi^{\prime}(s)=4p(1+s)\log(1+s)-4q(1-s)\log(1-s)-2[p(1+s)+q(1-s)]\log[p(1+s)^{2}+q(1-s)^{2}]-4s.$
Since $\psi^{\prime}(0)=0$, it is enough to prove $\psi^{\prime\prime}(s)\leq 0,s\in[-1,1]$. We continue to compute
$\psi^{\prime\prime}(s)=4p\log(1+s)+4q\log(1-s)-2\log[p(1+s)^{2}+q(1-s)^{2}]-\frac{4(s+p-q)^{2}}{p(1+s)^{2}+q(1-s)^{2}}.$
Therefore, the proof will be finished if for all $s\in[-1,1]$
$\phi(p):=4p\log(1+s)+4q\log(1-s)-2\log[p(1+s)^{2}+q(1-s)^{2}]\leq 0,\qquad p\in[0,1].$
To see this, note that $\phi(0)=\phi(1)=0$. So it remains to prove the convexity of $\phi$, which follows from
$\phi^{\prime}(p)=4\log(1+s)-4\log(1-s)-\frac{8s}{p(1+s)^{2}+q(1-s)^{2}}$
and
$\phi^{\prime\prime}(p)=\frac{32s^{2}}{[p(1+s)^{2}+q(1-s)^{2}]^{2}}\geq 0.$
Appendix B More on the local Bobkov inequality
Recall that in our proof for biased discrete hypercube case, we used [BG99, Theorem 3.2] stating that for probability spaces $(\Omega_{i},\nu_{i}),1\leq i\leq n$ and their product $(\Omega,\nu)$, we have
$\mathcal{I}(\mathbb{E}_{\nu}h)\leq\sqrt{2}\mathbb{E}_{\nu}\sqrt{\sumop\displaylimits_{i}\textnormal{Var}_{\nu_{i}}(h)}$ (46)
for any measurable $h:\Omega\to\{0,1\}$, where $\mathcal{I}$ is the Gaussian isoperimetric profile $\mathcal{I}:[0,1]\to[0,1/\sqrt{2\pi}]$. In the following, we explain that the local Bobkov inequality extends to more general semigroups.
Fix the probability spaces $(\Omega_{i},\mu_{i}),1\leq i\leq n$ and their product $(\Omega,\mu)$. Let $P_{t}^{(i)}$ denote the depolarizing Markov semigroup on $(\Omega_{i},\mu_{i})$, that is, $P_{t}^{(i)}$ has generator $L_{i}(f)=\mu_{i}(f)-f$. Denote by $P_{t}$ the product of $P_{t}^{(i)}$’s. We will show that for all $t\geq 0$, we have ($\Gamma$ denoting the carré du champs for $P_{t}$)
$\mathcal{I}(P_{t}h)\leq 2\sqrt{(1-e^{-t})}P_{t}\sqrt{\Gamma(h)},\qquad h:\Omega\to\{0,1\}.$ (47)
As before, we fix $t\geq 0$ and $x=(x_{1},\dots,x_{n})\in\Omega$. Consider the positive linear functional
$\varphi_{i}(g):=P_{t}^{(i)}(g)(x_{i}),\qquad g:\Omega_{i}\to\real,$
such that $\varphi_{i}(1)=1$. Again, by Riesz representation theorem, there exists a probability measure $\nu_{i}$ on $\Omega_{i}$ such that
$\mathbb{E}_{\nu_{i}}(g)=P_{t}^{(i)}(g)(x_{i})=e^{-t}g(x_{i})+(1-e^{-t})\mu_{i}(g),\qquad g:\Omega_{i}\to\real.$
Thus, for any $n\geq 1$ and $\nu:=\otimes_{i=1}^{n}\nu_{i}$ one has
$\mathbb{E}_{\nu}(g)=P_{t}(g)(x),\qquad g:\Omega\to\real.$
Applying (46) to the above $\nu_{i}$’s, one obtains
$\mathcal{I}(P_{t}h)(x)\leq\sqrt{2}\mathbb{E}_{\nu}\sqrt{\sumop\displaylimits_{i}\textnormal{Var}_{\nu_{i}}(h)(x)}.$
It remains to show
$\sumop\displaylimits_{i}\textnormal{Var}_{\nu_{i}}(h)(x)\leq 2(1-e^{-t})\Gamma(h)(x).$ (48)
To see this, note that if we denote by $\Gamma_{i}$ the carré du champs associated to $P_{t}^{(i)}$, then
$\Gamma_{i}(h)=\frac{1}{2}\left[L_{i}(h^{2})-2L_{i}(h)h\right]=\frac{1}{2}\left[\mu_{i}(h)-2\mu_{i}(h)h+h\right],\qquad h:\Omega_{i}\to\{0,1\}$
and
$\Gamma(h)=\frac{1}{2}\sumop\displaylimits_{1\leq i\leq n}\left[\mu_{i}(h)-2\mu_{i}(h)h+h\right],\qquad h:\Omega\to\{0,1\}.$
A direct computation shows that for all $h:\Omega\to\{0,1\}$
$\textnormal{Var}_{\nu_{i}}(h)=\mathbb{E}_{\nu_{i}}(h)-\mathbb{E}_{\nu_{i}}(h)^{2}$
$\qquad=(1-e^{-t})\left[e^{-t}(\mu_{i}(h)-h)^{2}+\mu_{i}(h)(1-\mu_{i}(h))\right]$
$\qquad\leq(1-e^{-t})\left[(\mu_{i}(h)-h)^{2}+\mu_{i}(h)(1-\mu_{i}(h))\right]$
$\qquad=(1-e^{-t})\left[\mu_{i}(h)-2\mu_{i}(h)h+h\right].$
All combined, we proved (48) and thus (47).